# Holmes: Towards Distributed Training Across Clusters with Heterogeneous NIC Environment

### v3

@@ Abstract
1 Introduction
@@
In recent years, the emergence of large language models (LLMs) [31, 39] such as GPT-3 [6], OPT [38], and LLaMA [32] has significantly transformed the landscape of natural language processing (NLP) by achieving state-of-the-art accuracy in a wide variety of tasks. Despite these remarkable advancements, LLM training is often accompanied by considerable costs, requiring the sustained operation of tens of thousands of advanced GPUs over extended periods. For instance, Meta’s 175 billion parameter OPT-175 was trained for 33 days on 1024 NVIDIA A100 GPUs. A slightly smaller model, LLaMa with 65 billion parameters, was trained for 21 days using a cluster of 2048 NVIDIA A100 GPUs [38, 32, 6, 37]. Conventionally, these models undergo training within specialized clusters that are equipped with homogeneous high-speed Remote Direct Memory Access (RDMA) network interface cards (NICs). However, the acquisition and maintenance of such dedicated clusters can present substantial financial and logistical challenges [32, 39].
@@
Figure 1: An overview of the Holmes framework. Holmes provides training support for a diverse range of LLM types and seamlessly integrates with contemporary mainstream LLM training frameworks. This flexibility extends to the training of LLMs in the heterogeneous NIC environment, where Holmes consistently achieves performance levels that are close to those achievable in the homogeneous NIC settings.
@@
RDMA technology has found extensive deployment in modern data centers [19, 10], offering low latency and high throughput benefits that are particularly advantageous for LLM training. RDMA technology primarily comprises two categories: InfiniBand and RoCE (RDMA over Converged Ethernet) [7, 12, 10]. Both of these technologies facilitate direct access to remote computer memory, bypassing the need for CPU intervention, thus leading to more efficient data transfer and communication. InfiniBand represents a dedicated networking technology, while RoCE is a protocol designed to implement RDMA within the Ethernet framework. Both InfiniBand and RoCE play an integral role in the construction of high-performance computing and data center networks [16, 12, 10]. Despite their widespread applications, the two networking technologies, InfiniBand and RoCE, are inherently incompatible with each other. This incompatibility poses challenges, especially in the context of training LLMs within or across clusters that incorporate various RDMA NICs. In such scenarios, the adoption of pure Ethernet for communication often emerges as the only viable option.
@@
Speeding up the LLM training process and minimizing associated costs have been persistent areas of research. Current LLM training frameworks, as evidenced in recent surveys [39], which feature prominent models like Megatron-LM [30, 23], DeepSpeed [26, 31, 1], and FairScale [8], predominantly prioritize speed within homogeneous data center networks. Consequently, these frameworks may not be easily adaptable to the diverse landscape of the heterogeneous RDMA NIC environment. To address these challenges, there is a need for an open and unified LLM training framework that can seamlessly accommodate the heterogeneity of RDMA NICs across various clusters, thus alleviating the operational complexity associated with large-scale data center networks.
@@
Specifically, our framework tracks hardware topology and network characteristics, dynamically assigning parallel work groups to suitable GPU devices during LLM training. The internal scheduling method, which is applicable to the heterogeneous NIC environment, is complemented by well-designed optimizations in both data and model parallelism [35, 14, 21, 20]. As a result, Holmes can effectively harness the full potential of the real-world production infrastructure. This includes the use of high-speed RDMA NICs for data parallelism, which exhibits a high communication overhead [20, 25, 28], and Ethernet NICs for pipeline model parallelism, known for its lower communication overhead [14]. We demonstrate impressive performance in using our framework to train LLM across clusters equipped with heterogeneous NICs.
@@ 2 Problem Formulation
In this section, we first provide an overview of the background and motivations behind our work. Following this, we introduce different NIC environments. Finally, we describe the problem addressed in this paper and proceed to formalize it.
2.1 Background and Motivation
2.1.1 Communication Technologies
In addition to Ethernet, which is a widely used protocol for network connection, RDMA, NVLink, and PCI-E are three distinct technologies or protocols used to support varying degrees of data transmission and communication [9, 17, 10].
  • RDMA is a network technology enabling direct memory access between computer systems without CPU involvement in data transfer. It achieves high-performance remote memory access by bypassing the CPU, reducing latency.
  • NVLink is a high-speed technology connecting multiple GPUs for enhanced parallel computing. It provides a high-bandwidth, low-latency communication channel with direct memory access, improving data transfer efficiency in multi-GPU systems.
  • PCI-E is a high-speed serial bus standard connecting diverse hardware in computers, including graphics cards, storage, and network devices. It offers a versatile, high-bandwidth, and low-latency data transfer channel through point-to-point connections. While PCI-E is commonly used for various hardware, it may have limitations in data transfer efficiency compared to technologies like NVLink, specifically designed for GPU-to-GPU communication..
In summary, RDMA enables high-performance network communication, NVLink enhances communication between GPUs, and PCI-E is a versatile and widely used internal computer bus standard. These technologies are crucial in various domains and applications, particularly for distributed training.
Figure 2: An illustration of the scheduling problem in heterogeneous NIC environments, with the objective of minimizing communication costs for distributed LLM training tasks.
2.1.2 LLM Infrastructure
@@
Training LLMs requires an advanced infrastructure, including GPU clusters, high-speed RDMA NICs, and ample storage capacity. In particular, the demand for NVIDIA A100 GPU devices often exceeds 5,000 [23]. As LLM continues to grow in size, this demand is expected to increase further. However, the development of cluster infrastructure has lagged behind rapid advances in LLM. Many existing GPU clusters, constructed several years ago, lack the capacity to accommodate more than 1,000 NVIDIA A100 GPU devices, and network connectivity between GPU clusters in different locations is often limited. Consequently, the need to reconstruct larger GPU clusters for LLM research has become a necessity.
@@
Moreover, the RDMA technology has two implementations: InfiniBand and RoCE. Different GPU clusters may be equipped with different types of RDMA NIC, and these two variants are not mutually compatible. Existing training frameworks, such as Megatron-LM [30, 23] and Megatron-DeepSpeed [26, 1], face challenges in training LLM in the heterogeneous NIC environment. Consequently, they are unable to fully harness GPU devices in the heterogeneous clusters. This limitation leads to a suboptimal utilization of GPU computing resources.
2.1.3 Motivation
These practical challenges in infrastructure prompt us to reconsider LLM training, focusing on the effective utilization of GPU devices and RDMA NICs across multiple clusters. Our objective is to develop a unified LLM training framework with the following key goals:
  • Compatibility: The training framework should seamlessly operate with different RDMA NICs across multiple GPU clusters. To accomplish this, we redesign existing schemes for parallelism methods.
  • Scalability: The training framework should easily scale to accommodate thousands of GPU devices. For this purpose, we have developed Holmes based on the Megatron-LM framework, a popular LLM training framework known for successfully handling training iterations on models with up to a trillion parameters.
  • Availability: Modern large-scale GPU clusters are consistently equipped with RDMA NICs. Our training framework maximizes the use of these available machine resources while ensuring high performance.
2.2 NIC Environments
@@
We consider the scenario involving a group of GPU devices engaged in collaborative LLM training. These GPUs are from clusters equipped with different types of RDMA NIC, including InfiniBand or RoCE. We summarize the following two cases within this context.
@@ 2.3 Problem Description
  • TFLOPS stands for the achieved teraFLOP/s per GPU, representing the number of floating-point operations a GPU can perform in one second. It measures the GPU utilization rate, and the computational formula aligns with that in [23]. Further details are provided in the Experiment section (see Equation 6).
  • Throughout stands for the number of samples processed per second during the LLM training procedure. It measures the end-to-end training speed.
2.4 Formalization
@@ 3 Holmes Design
Figure 3: An illustration of parallelism mechanism in Holmes. In this case, a transformer-based model with 6 layers is trained across 2 clusters. Each cluster comprises 2 nodes, each equipped with 4 GPU devices. Communication between Node 1 and Node 2 utilizes InfiniBand, while Node 3 and Node 4 use RoCE. However, there is no high-speed interconnect between the two clusters, and communication between them relies solely on low-speed Ethernet. For parallelism settings, the degrees of data, tensor, and pipeline parallelisms are 2, 2, and 4, respectively. Pipeline parallelism is implemented between the two clusters using Ethernet. The model’s layers are unevenly partitioned into 2 stages and further distributed to different GPU devices. Data parallelism is performed within each cluster using RDMA, and tensor parallelism is implemented within each node using NVLink.
@@
In this section, we present the Holmes design. Holmes is developed within the framework of Megatron-LM [30, 23], a popular tool for LLM training. Our approach re-optimizes two types of parallelism techniques: model parallelism and data parallelism. Model parallelism is further categorized into tensor parallelism and pipeline parallelism.
In distributed training, pipeline parallelism typically has less communication overhead than data parallelism because different GPU devices handle sequential computational stages with less inter-device communication between them [14, 21]. For data parallelism, data is divided into segments, and each device independently processes a portion. These units simultaneously perform identical operations on separate data, leading to increased communication for synchronization and greater overhead [26, 28]. Holmes enhances the built-in communication methods for pipeline and data parallelism in Megatron-LM. Figure 3 illustrates the parallelism mechanism in Holmes. Further details will be discussed in the subsequent parts.
3.1 Model Parallelism
3.1.1 Tensor Model Parallelism
@@
Tensor model parallelism involves distributing individual layers of the model across multiple GPU devices. In the tensor parallelism of our framework, different layers of the model are assigned to various GPU devices within the same node, enabling concurrent processing of these layers, and thereby accelerating the training process. As a result, the communication operations associated with tensor parallelism predominantly take place within the same node, leveraging technologies such as PCI-E [15, 17] or NVLink [9] to facilitate inter-device communication.
Holmes adopts the same partitioning strategy used by Megatron-LM for transformer layers, which serve as the fundamental components of LLMs [23]. Following the partitioning process, we obtain the tensor parallel matrix denoted as $[\textbf{TP}]\in\mathbb{Z}_{+}^{(p\ \cdot\ d)\times t}$. This matrix contains various parallel groups, and the GPU devices within each group belong to the same node. We represent $[\textbf{TP}]_{i,j}$ as the $j$-th device in the $i$-th tensor parallel group, which can be formalized as follows:
${[\textbf{TP}]_{i,j}=\operatorname{rank}_{(i-1)\cdot t+j}}\quad_{j\in\{1,2\ ,...,\ t\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ d\}}$ (1)
Due to space limitations, we omit the further technical introduction in this section. For more detailed information on tensor parallelism, refer to [23].
3.1.2 Pipeline Model Parallelism
In traditional pipeline parallelism, the model layers are segmented into different computational stages, each assigned to different GPU devices. These devices execute the stages sequentially, transferring their results to subsequent stages. Communication operations occur through the exchange of data between these devices. When dealing with models that contain repeated transformer blocks, each GPU device can effectively manage an equal number of these blocks. This process involves breaking a batch into smaller micro-batches and executing them in a pipeline fashion [14, 21, 22]. We use periodic pipeline flushes to maintain the synchronization of optimizer steps between GPU devices.
The implementation of our pipeline parallelism is similar to PipeDream-Flush [22]. To support LLM training across clusters equipped with heterogeneous NICs, we make the following innovative improvements.
Cross-Cluster Pipeline Parallelism.
@@ 3.1.2 Pipeline Model Parallelism
Current LLM training frameworks exclusively support LLM training with low-speed Ethernet NICs within clusters that possess heterogeneous NICs. To address these limitations, we have introduced a solution to enable LLM training in the heterogeneous NIC environment. Capitalizing on the relatively low communication cost of pipeline parallelism, we establish pipeline parallelism groups among multiple clusters that lack high-speed interconnects. Consequently, these pipeline parallelism groups consist of GPU devices connected via heterogeneous NICs, and their sole means of communication is through Ethernet due to NIC incompatibility. Conversely, owing to the division of pipeline parallelism groups, the data parallelism groups can comprise GPU devices connected via homogeneous NICs, enabling communication via RDMA NICs. In terms of tensor parallelism, it is confined to the same node and communicates via PCI-E or NVLink interfaces.
NIC Env TFLOPS Throughtout Bandwidth (Gbps)
@@
Table 1: Comparison of TFLOPS and Throughput metrics when training a GPT model with 3.6 billion parameters on 4 nodes using InfiniBand, RoCE, and Ethernet respectively. Each node is equipped with 8 NVIDIA A100 GPUs. We also list the bandwidth of different NICs in the 3rd column.
Considering that data parallelism relies more on high-speed RDMA NICs than pipeline parallelism does, our Cross-Cluster Pipeline Parallelism method significantly enhances LLM training performance by scheduling pipeline parallelism groups across different clusters in heterogeneous NIC environments.
Self-Adapting Pipeline Partition.
Traditional pipeline parallelism was initially designed for training LLM in the homogeneous NIC environment. In such settings, the model layers are uniformly divided into different stages, ensuring relatively equal computational speeds across various stages of pipeline parallelism. This uniform partitioning is considered the most suitable choice, as all GPU devices exhibit equal computational speeds [23]. However, in the context of a heterogeneous NIC environment, different GPU devices may demonstrate varying computational speeds due to the diverse NICs to which they are connected. Table 1 presents a comparison of GPU performance using three different types of NIC during LLM training. Even if InfiniBand and RoCE NICs have the same bandwidth, the GPU device equipped with different types of NIC may exhibit significant variations in actual computational speed (TFLOPS). The observed disparities in TFLOPS and Throughput metrics underscore the inadequacy of a straightforward uniform partition for pipeline parallelism in the heterogeneous NIC environment.
$N_{ib}=\lfloor\frac{\alpha\cdot S(\textsc{IB})}{S(\textsc{IB})+S(\textsc{RoCE})}\cdot N\rfloor$ (2)
where $\alpha$ is the hyper-parameter. Similarly, the calculation of $N_{roce}$ follows the same method, ensuring $N_{ib}+N_{roce}=N$. It can be easily extended to scenarios where the pipeline parallel size is more than 2. Our objective is to allocate a greater number of model layers to the GPU device connected to the faster NIC.
Summary.
We introduce two innovative enhancements to pipeline parallelism: optimizing communication operations and adapting to the heterogeneous NIC environment. We illustrate the process with a case presented in Figure 3.

### v4

@@ Abstract.
Large language model, Training framework, Heterogeneous NIC environment
* The corresponding author.
ccs: Computer systems organization Distributed architectures
1. Introduction
@@
In recent years, the emergence of large language models (LLMs) (Smith et al., 2022; Zhao et al., 2023) such as GPT-3 (Brown et al., 2020), OPT (Zhang et al., 2022), and LLaMA (Touvron et al., 2023) has significantly transformed the landscape of natural language processing (NLP) by achieving state-of-the-art accuracy in a wide variety of tasks. Despite these remarkable advancements, LLM training is often accompanied by considerable costs, requiring the sustained operation of tens of thousands of advanced GPUs over extended periods. For instance, Meta’s 175 billion parameter OPT-175 was trained for 33 days on 1024 NVIDIA A100 GPUs. A slightly smaller model, LLaMa with 65 billion parameters, was trained for 21 days using a cluster of 2048 NVIDIA A100 GPUs (Zhang et al., 2022; Touvron et al., 2023; Brown et al., 2020; Yuan et al., 2022). Conventionally, these models undergo training within specialized clusters that are equipped with homogeneous high-speed Remote Direct Memory Access (RDMA) network interface cards (NICs). However, the acquisition and maintenance of such dedicated clusters can present substantial financial and logistical challenges (Touvron et al., 2023; Zhao et al., 2023).
@@
Figure 1. An overview of the Holmes framework. Holmes provides training support for a diverse range of LLM types and seamlessly integrates with contemporary mainstream LLM training frameworks. This flexibility extends to the training of LLMs in the heterogeneous NIC environment, where Holmes consistently achieves performance levels that are close to those achievable in the homogeneous NIC settings.
@@
RDMA technology has found extensive deployment in modern data centers (Li et al., 2023; Gao et al., 2021), offering low latency and high throughput benefits that are particularly advantageous for LLM training. RDMA technology primarily comprises two categories: InfiniBand and RDMA over Converged Ethernet (RoCE) (Chen et al., 2016; Guo et al., 2016; Gao et al., 2021). Both of these technologies facilitate direct access to remote computer memory, bypassing the need for CPU intervention, thus leading to more efficient data transfer and communication. InfiniBand represents a dedicated networking technology, while RoCE is a protocol designed to implement RDMA within the Ethernet framework. Both InfiniBand and RoCE play an integral role in the construction of high-performance computing and data center networks (Kalia et al., 2016; Guo et al., 2016; Gao et al., 2021). Despite their widespread applications, the two networking technologies, InfiniBand and RoCE, are inherently incompatible with each other. This incompatibility poses challenges, especially in the context of training LLMs within or across clusters that incorporate various RDMA NICs. In such scenarios, the adoption of pure Ethernet for communication often emerges as the only viable option.
@@
Speeding up the LLM training process and minimizing associated costs have been persistent areas of research. Current LLM training frameworks, as evidenced in recent surveys (Zhao et al., 2023), which feature prominent models like Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b), DeepSpeed (Rasley et al., 2020; Smith et al., 2022; MicroSoft, 2023), and FairScale (FairScale authors, 2021), predominantly prioritize speed within homogeneous data center networks. Consequently, these frameworks may not be easily adaptable to the diverse landscape of the heterogeneous RDMA NIC environment. To address these challenges, there is a need for an open and unified LLM training framework that can seamlessly accommodate the heterogeneity of RDMA NICs across various clusters, thus alleviating the operational complexity associated with large-scale data center networks.
@@
Specifically, our framework tracks hardware topology and network characteristics, dynamically assigning parallel work groups to suitable GPU devices during LLM training. The internal scheduling method, which is applicable to the heterogeneous NIC environment, is complemented by well-designed optimizations in both data and model parallelism (Xing et al., 2015; Huang et al., 2019; Narayanan et al., 2019; Li et al., 2020). As a result, Holmes can effectively harness the full potential of the real-world production infrastructure. This includes the use of high-speed RDMA NICs for data parallelism, which exhibits a high communication overhead (Li et al., 2020; Rajbhandari et al., 2020; Ren et al., 2021), and Ethernet NICs for pipeline model parallelism, known for its lower communication overhead (Huang et al., 2019). We demonstrate impressive performance in using our framework to train LLM across clusters equipped with heterogeneous NICs.
@@
Figure 2. An illustration of the scheduling problem in heterogeneous NIC environments, with the objective of minimizing communication costs for distributed LLM training tasks.
@@ 2. Preliminary
In this section, we first provide an overview of LLM infrastructure. Following this, we introduce different NIC environments. Finally, we describe and formalize the problem addressed in this paper.
2.1. LLM Infrastructure
@@
Training LLMs requires an advanced infrastructure, including GPU clusters, high-speed RDMA NICs, and ample storage capacity. In particular, the demand for NVIDIA A100 GPU devices often exceeds 5,000 (Narayanan et al., 2021b). As LLM continues to grow in size, this demand is expected to increase further. However, the development of cluster infrastructure has lagged behind rapid advances in LLM. Many existing GPU clusters, constructed several years ago, lack the capacity to accommodate more than 1,000 NVIDIA A100 GPU devices, and network connectivity between GPU clusters in different locations is often limited. Consequently, the need to reconstruct larger GPU clusters for LLM research has become a necessity.
@@
Moreover, the RDMA technology has two implementations: InfiniBand (IB) and RDMA over Converged Ethernet (RoCE). Different GPU clusters may be equipped with different types of RDMA NIC, and these two variants are not mutually compatible. Existing training frameworks, such as Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b) and Megatron-DeepSpeed (Rasley et al., 2020; MicroSoft, 2023), face challenges in training LLM in the heterogeneous NIC environment. Consequently, they are unable to fully harness GPU devices in the heterogeneous clusters. This limitation leads to a suboptimal utilization of GPU computing resources.
2.2. NIC Environments
@@
We consider the scenario involving a group of GPU devices engaged in collaborative LLM training. These GPUs are from clusters equipped with different types of RDMA NIC, including IB or RoCE. We summarize the following two cases within this context.
@@ 2.3. Problem Description
  • TFLOPS stands for the achieved teraFLOP/s per GPU, representing the number of floating-point operations a GPU can perform in one second. It measures the GPU utilization rate, and the computational formula aligns with that in (Narayanan et al., 2021b).
  • Throughput stands for the number of samples processed per second $(samples/s)$ during the LLM training procedure. It measures the end-to-end training speed.
2.4. Formalization
@@ 3. Holmes Design
Figure 3. An illustration of parallelism mechanism in Holmes. In this case, a transformer-based model with 6 layers is trained across 2 clusters. Each cluster comprises 2 nodes, each equipped with 4 GPU devices. Communication between Node 1 and Node 2 utilizes InfiniBand, while Node 3 and Node 4 use RoCE. However, there is no high-speed interconnect between the two clusters, and communication between them relies solely on low-speed Ethernet. For parallelism settings, the degrees of data, tensor, and pipeline parallelisms are 2, 2, and 4, respectively. Pipeline parallelism is implemented between the two clusters using Ethernet. The model’s layers are unevenly partitioned into 2 stages and further distributed to different GPU devices. Data parallelism is performed within each cluster using RDMA, and tensor parallelism is implemented within each node using NVLink.
@@
In this section, we present the Holmes design. Holmes is developed within the framework of Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b), a popular tool for LLM training. Our approach re-optimizes two types of parallelism techniques: model parallelism and data parallelism. Model parallelism can be further categorized into tensor parallelism and pipeline parallelism. Holmes enhances the built-in communication methods for pipeline and data parallelism in Megatron-LM through optimizations including Cross-cluster Pipeline Parallelism, Automatic NIC Selection and Self-Adapting Pipeline Partition. Figure 3 illustrates an overall design of the parallelism mechanism in Holmes. Further details will be discussed in the subsequent parts.
3.1. Cross-cluster Pipeline Parallelism
3.1.1. Parallelism Strategy Design
In order to facilitate a large-scale training job across multiple clusters, a dedicated parallelism strategy is required. As depicted in Figure 3, the basic parallelism techniques could be categorized into data parallelism, pipeline parallelism, and tensor parallelism. Each parallelism technique utilizes different communication patterns and therefore exhibits a specific network requirement. Hence, the strategies of various parallelism techniques should be coherent to the heterogeneous NIC environments.
For tensor parallelism, it involves distributing individual layers of the model across multiple GPU devices. The computation paragidm within tensor parallelism is the operation on splited model weights and complete activation tensors. Substantial broadcast and gather operations on activation tensors are therefore implemented, resulting in substantial overhead in communication.
@@ 3.1.1. Parallelism Strategy Design
In the tensor parallelism of our framework, different parts of the model’s layers are assigned to various GPU devices within the same node, enabling concurrent processing of these layers, and thereby accelerating the training process. As a result, the communication operations associated with tensor parallelism predominantly take place within the same node, leveraging technologies such as PCI-E (Kaldewey et al., 2012; Li et al., 2019) or NVLink (Foley and Danskin, 2017) to facilitate inter-device communication. Holmes adopts the same partitioning strategy used by Megatron-LM for transformer layers, which serve as the fundamental components of LLMs (Narayanan et al., 2021b).
For pipeline parallelism, the model layers are segmented into different computational stages, each assigned to different GPU devices. These devices execute the stages sequentially, transferring their results to subsequent stages. Send and receive communication operations occur through the exchange of data between these devices. The communication workload of pipeline parallelism is comparably low among the three parallelism techniques. Moreover, it is possible to overlap the computation and communication between parallel stages. Therefore, the network bottleneck between clusters could be concealed by carefully designed pipeline parallelism strategy.
@@
The implementation of our pipeline parallelism is similar to PipeDream-Flush (Narayanan et al., 2021a). We use periodic pipeline flushes to maintain the synchronization of optimizer steps between GPU devices. Current LLM training frameworks exclusively support LLM training with low-speed Ethernet NICs within clusters that possess heterogeneous NICs. To address the associated limitations, we have developed a solution to facilitate LLM training in heterogeneous NIC environment. Capitalizing on the relatively low communication cost of pipeline parallelism than tensor parallelism and data parallelism, we establish pipeline parallelism groups among multiple clusters that lack high-speed interconnects. Consequently, these pipeline parallelism groups consist of GPU devices connected via heterogeneous NICs, and their sole means of communication is through Ethernet due to NIC incompatibility. Conversely, owing to the division of pipeline parallelism groups, the data parallelism groups can comprise GPU devices connected via homogeneous NICs, enabling communication via RDMA NICs. In terms of tensor parallelism, it is confined to the same node and communicates via PCI-E or NVLink interfaces.
NIC Env TFLOPS Throughtput Bandwidth (Gbps)
@@
Table 1. Comparison of TFLOPS and Throughput metrics when training a GPT model with 3.6 billion parameters on 4 nodes using InfiniBand, RoCE, and Ethernet respectively. Each node is equipped with 8 NVIDIA A100 GPUs. We also list the bandwidth of different NICs in the 3rd column.
For data parallelism, each GPU device possesses a complete replica of the model, while the input dataset is partitioned among these GPU devices. As previously noted, the communication cost associated with data parallelism is considerable. Communication operations in data parallelism occur in the steps including data partition, gradient computation, gradient aggregation and model update. Communication between GPUs primarily occurs during the gradient aggregation step using all-reduce operations, facilitated by high-speed RDMA NIC and libraries like NCCL (NVIDIA, 2021), ensuring the synchronization of gradients between GPUs and maintaining consistent model parameters across all devices.
Similarly to pipeline parallelism, traditional data parallelism implementation does not take into account the heterogeneous NIC environments, and it can only use the low-speed Ethernet NIC when training LLMs in the heterogeneous environment. To address the problem and accelerate the data parallelism in both homogeneous and heterogeneous environment, we introduce the following enhancement to data parallelism: an optimal selection strategy of NICs, which helps creating data parallelism groups based on the actual network environment, ensuring that GPU devices within the same data parallelism group are connected to the same type of NIC, while different data parallelism groups can select different NICs. Detailed implementation is presented in the subsequent parts. Considering that data parallelism relies more on high-speed RDMA NICs than pipeline parallelism does, our Cross-Cluster Pipeline Parallelism method has potential to significantly improve LLM training performance by scheduling pipeline parallelism groups across different clusters in heterogeneous NIC environments.
3.1.2. Formalization
In order to implement the aforementioned design, we propose a formalized specification of the cross-cluster pipeline parallelism strategy in terms of parallel matrices over various parallel groups of devices.
Following the partitioning process, we represent $[\textbf{TP}]_{i,j}$ as the $j$-th device in the $i$-th tensor parallel group, which can be formalized as follows:
(1) ${[\textbf{TP}]_{i,j}=\operatorname{rank}_{(i-1)\cdot t+j}}\quad_{j\in\{1,2\ ,...,\ t\}}^{i\in\{1,2\ ,...,\ p\ \cdot\ d\}}$
For the $i$-th tensor parallel group $[\textbf{TP}]_{i,*}$, it comprises of $t$ devices from the same node.